Sandy Bank, Inc., makes one model of wooden canoe. And, the infomation for it follows: Number of canoes produced and sold Total costs 700 S 92,500 S129 500 $157.250 S271,000 S308,000 $335,750 S 185.00 S18500 185.00 S512.00 14000 $ 395.00 500 850 ariable costs Fixed costs Total costs Cost per unit ariable cost per unit Fixed cost per unit Total cost per unit 357.00 255.00 210.00 Required: Suppo e tha andy ank re ses s seling price to S500 per c noe alculele its new break e ven port n unts 테nd in ales do lear Do no round intermediate calculations Round your final answers o nearest whole number. New Break Even Units Cances Break-Even Sales Revenue 2. If Sandy Bank sells 1,560 canoes, compute its margin of safety in dollars arnd as a percentage of sales (Use the new sales price of S500 ) (Round your answers to the nearest whole number.) Margin of Safety Percentage of Sales 3. Calculate the number of canoes that Sandy Bank must sell at $500 each to generate S120,000 proft (Round your answer to the nearest whole number.) arget S Canoes
Answer #1

1.New Selling Price per Unit = $500

Less: Variable Cost per Unit = $185

Contribution Margin per Unit = $315

Break Even Units = Total Fixed Costs/Contribution Margin per Unit

= 178,500/315

= 567 units

Break even sales revenue = Break even sales units*Selling Price per Unit

= 567*500

= $283,500

2.Sandy Bank Sells 1,560 canoes

Margin of Safety in Units = 1,560 – 567 = 993 units

In Dollars = 993*500 = $496,500

% of Sales = 993/1560 = 64%

3.Desired Profit = $120,000

+ Fixed costs = $178,500

Desired Contribution Margin = $298,500

Contribution Margin per Unit = $315

Number of canoes required to be sold = 298,500/315

= 948 canoes